# Intra-night optical variability of peculiar narrow-line Seyfert 1 galaxies with enigmatic jet behavior

### 08 Jan 2024 v1

@@ Acknowledgements
We are thankful to the anonymous referee for providing comments and suggestions, which helped us to improve the manuscript considerably. We thank Dr. Mudit Srivastava and Dr. Vipin Kumar for their help in carrying out observations with MFOSC-P on the 1.2m telescope and also acknowledge the help of Dr. Vishal Joshi for using FOC on the 2.5m telescope. We also thank the concerned ARIES and PRL staff for assistance during the observations obtained at the 1.04m Sampurnanand Telescope (ST), 1.30m Devasthal Fast Optical Telescope (DFOT) of ARIES and 1.20m, 2.50m Mt. Abu telescopes of PRL, which are the national facilities run and managed by ARIES and PRL, respectively.

### 20 Jan 2024 v2

@@ Acknowledgements
We acknowledge the Director of Physical Research Laboratory (PRL) and the Department of Space (DOS), Government of India, for supporting this research work and VO’s postdoctoral fellowship. The PRL runs and supports the Observational activities with the 1.2m and 2.5m telescopes at Mt. Abu. We acknowledge the PRL Mt. Abu Observatory staff for our observations with the PRL’s 1.2m and 2.5m telescopes. This work has used observations from the Mt. Abu Faint Object Spectrograph and Camera-Pathfinder (MFOSC-P) and we acknowledge the help from Dr. Vipin Kumar and the entire MFOSC-P team led by Dr. Mudit K. Srivastava. This work has also used observations with the Faint Object Camera (FOC) on the 2.5m telescope and we acknowledge the efforts of the FOC development team led by Prof. Abhijit Chakraborty and Dr. Vishal Joshi. We also thank the Aryabhatta Research Institute of Observational Sciences (ARIES) authorities and staff for their assistance during the observations taken from the 1.04m Sampurnanand Telescope (ST) at Nainital, 1.30m Devasthal Fast Optical Telescope (DFOT) at Devasthal, both ST and DFOT are run by ARIES. We thank the anonymous referee for providing useful comments and suggestions, which helped us to improve the manuscript.